# An open framework for archival, reproducible, and transparent science

### v1

@@
Sabar Dasgupta Department of Bioengineering Paul Nuyujukian Department of Bioengineering Department of Neurosurgery Department of Electrical Engineering Wu Tsai Neurosciences Institute Stanford Bio-X
@@ 1 Abstract
Digital computational outputs are now ubiquitous in the research workflow and the way in which these data are stored and cataloged is becoming more standardized across fields of research. However, even with accessible data and code, the barrier to recreating figures and reproducing scientific findings remains high. What is generally missing is the computational environment and associated pipelines in which the data and code are executed to generate figures. The archival, reproducible, and transparent science (ARTS) open framework incorporates containers, version control systems, and persistent archives through which all data, code, and figures related to a research project can be stored together, easily recreated, and serve as an accessible platform for long-term sharing and validation. If the underlying principles behind this framework are broadly adopted, it will improve the reproducibility and transparency of research.
@@ 2 Introduction
Over the past few decades, researchers have recognized the large effort required to replicate experiments and the rate at which these replication studies produce negative results [28]. Dubbed a “replicability crisis” [43], the current paradigm for communicating experimental methods and distributing research outputs impedes future collaboration and the development of public trust in findings. To improve replicability, both the processes for data collection and those for creating reproducible analyses must be improved. Reproducibility is the primary focus of this manuscript with a goal of minimizing the time, effort, and expertise required to analyze data from a study beyond its published manuscript [25].
@@
By the end of of 2025, research funded by the US federal government must make data accessible at the time of publishing [54, 49]. Each funding agency has their own specific policies for data sharing [37, 40, 50, 32, 13, 51]. While these policies offer guidance for researchers pursuing collaboration and open science, they do not outline how data, and more importantly, entire workflows should be shared and made accessible. The details around implementing computational best practices are left up to researchers, which is generally outside their scope of expertise and takes the focus away from core scientific work.
There are a number of standards put out by the scientific community to address these concerns ranging from generic (and not necessarily research-specific) to highly focused on one field of research. Data standards include the FAIR Guiding Principles [56] and DataFed [12]. While FAIR is a good set of principles surrounding the sharing of data, it alone does not assure reproducibility as it does not consider computation. Standards for reproducibility must advise on not only how data is stored, but also how it is manipulated and presented in analysis.
@@
Frameworks such as Spyglass (for neuroscience) or Galaxy (for bioinformatics) recommend field-specific workflows with associated software and are an excellent step towards integrating various software frameworks, but can be fragile as they rely on the permanence and compatibility of external services and APIs [30, 24]. Journals are starting to place a focus on software quality and availability and are recommending ways in which scientific code should be written in order to ensure accessibility [29]. Major journals such as Nature Scientific Data also have specific requirements and recommendations for deposition of data and code, but fall short of promoting accessible pipeline reproducibility [45]. Specifically, Nature and Science now prohibit the use of GitHub as a code deposition framework because they can be altered at a later time [35, 44] and Cell strongly recommends archives that support digital longevity [4]. New SaaS tools such as Code Ocean have been developed to address these issues and offer reproducibility and analysis pipeline visualization as a service [7]. Code Ocean uses Amazon as a cloud provider to reproduce analyses and allows exporting of the files that define the workflow environment—Containerfiles—so that they can be run elsewhere. While this is a useful tool for increasing reproducibility for the journal review process [5], it does not define an open standard and is a centralized, commercial solution that is not accessible during the research lifecycle. The calkit Python package is another useful reproducibility tool and acts as a wrapper around some of the needed technologies discussed later in this manuscript. [3]. These existing environment solutions do not yet define an open framework that aims to minimize compliance burden and the time and effort required to reproduce a research workflow many years down the line. Neither highly opinionated field-specific frameworks nor existing generic research workflow aids provide end-to-end support at the computational environment level for a wide array of fields.
@@
The medal scale (pertaining to life sciences) [25] recommends automating dependency installation and describes a gold standard where analyses are reproducible with a signal command. Unless these dependencies are versioned and running in a consistent environment, there is no guarantee that they will produce deterministic outputs or even run without error in the future. While reproduction effort is a crucial metric, a further key consideration is how far into the future the workflow is able to be run in one step. To truly meet this gold standard the complete software pipelines associated with experiment analyses must be captured as a container image—a binary file that bundles all dependencies, including the operating system, needed to recreate the experiment outputs.
@@
Until recently, reproducing pipelines encompassing a workflow was difficult, if not impossible, without locking in to specific computational platforms. Researchers concerned with reproducibility will often specify versioned dependencies and the operating system they used to run their code, but this may miss hidden dependencies, and further burdens researchers attempting to reproduce findings with the task of recreating the computational environment. For example, code run in a Python virtual environment installed with the same pinned dependencies may produce different outputs depending on the operating system version or the versions of other packages installed on the operating system. Further, these approaches assume that software library repositories such as PyPI for for Python or CRAN for R will continue to exist decades into the future. Containers, now “a pillar of scientific computing”, offer a solution to environment management that allows checkpointing entire workflows while being relatively operating system agnostic across local and cloud settings [38]. This manuscript describes a framework which aims to be a flexible blueprint that is compatible with as many existing systems as possible. Specific recommendations are made that allow pipeline reproducibility and define a compatibility layer for those looking to verify results and collaborate.
3 Framework Definition
@@
The Archival, Reproducible, and Transparent Science (ARTS) open framework for accessibly sharing research involving data collection and computational analyses is defined here. The goal of this framework is to provide those performing computational analyses that produce research outputs (digital artifacts) a set of best practices for tracking the evolution of a computational research effort, archiving data and results, sharing said artifacts, and making it simple for others to reproduce results and work within a consistent analysis environment. The framework can be used collaboratively, incorporating edits and sharing across multiple individuals and groups. Pieces of the ARTS open framework may also be omitted to define best practices for partial workflows—for example, archiving just datasets, archiving code that operates on already-archived datasets, or archiving a standalone computational environment that is used in many experiments.
@@
  • Data and code provenance documented using trusted timestamping
@@
The points outlined above and described in detail below should ideally be incorporated early into the research lifecycle. That way, when it is time to create a deposition and archive the research output, adoption of the framework does not pose an undue burden on the submitter. Further, having a portable deposition that can be uploaded to multiple archives makes research outputs more resilient.
@@ 3.1 Archives
One goal of archiving experimental results is to make all aspects of analysis available to future experimenters for decades to come: processing data, running models, and generating figures. In order to achieve this, it is vital that the archive service housing data will likely be in existence and freely accessible as far as possible into the future, or that depositions are able to be archived in a decentralized manner. When choosing an archive, there a few key considerations to make:
An ideal archive must be publicly accessible, meaning that repositories are available for download without paywalls or logins [45]. Accessible archives must also support open licenses and allow works to be searchable by included metadata like keywords. In the ideal case, an accessible repository would enable each individual file in a deposition to be individually addressable via a static URL, with a common URL prefix for the entire deposition, and the file contents accessible under the same name and directory structure as uploaded by the researcher.
@@
Archives must allow for persistent storage of artifacts. Once results are published, they should remain immutable and allowances for future updates or edits must result in an immutable history that includes the original publication. If history is not immutable, trusted timestamping information must be provided as described below. Persistent archives also must support persistent identifiers of some kind. Ideally this should be a digital object identifier (DOI), but a persistent URL (PURL) that is not a DOI may suffice if the service providing the DOI does not support other aspects of the framework. In these cases, associated articles should be published with a DOI and include a PURL that references the packaged experiment artifacts. Persistent archive services should also store a copy of data across multiple geographic regions so that they are resilient to external events that might result in data loss such as fires or flooding.
@@
An ideal archive should also be trusted. Requirements for trust cannot be singularly defined in a broad sense, but at a minimum, an archive service should be established and well supported with long-term financial support and a dedicated community. Archives backed by nonprofit institutions are also preferable as they are less likely to change their policies due to economic pressures. Providing for an optional confidential review process of artifacts also promotes trust in the archive and the data it houses [45].
@@
Figure 2: No popular research archive currently supports all aspects of the ARTS open framework, but Zenodo supports all aspects outside of hosting a container registry and allowing direct git clones of analysis code . Each column (besides storage limits) defines a question with checks and X’s indicating yes and no, respectively. Columns definitions are as follows: 3rd party upload: Does the archive allow anyone to upload data? Immutable: Once published, is deposition content locked from edits? DOI: Does the archive supply a Digital Object Identifier for uploads? PURL: Does the archive supply a persistent URL for uploads? Geographic redundancy: Is data uploaded to the archive backed up in separate geographic locations? Reviews: Does the archive provide a process for privately and confidentially sharing uploads to reviewers before publishing? Nonprofit: Does the maintainer of the archive operate under nonprofit status? Storage limits: Maximum file and deposition size as well as any other deposition limits. Bare repo: Does the archive allow depositions to be directly git cloned? Container pull: Does the archive implement a container registry to allow images to be pulled directly by container engines?
@@
Some academic institutions offer in-house archive solutions: the Dataverse Project originally developed at Harvard, DSpace developed between MIT and HP, and the Stanford Digital Repository (SDR) are some examples of archives that promote open access and provide persistent identifiers. DSpace and Dataverse are open-source tools with instances hosted by many partner organizations while the SDR is only available to Stanford affiliates. There are also popular and journal-recommended third-party solutions such as Zenodo, Dryad, or figshare if a field-specific data repository cannot be found. Dryad allows up to 300GB of storage per submission with a publication fee, while Zenodo and figshare do not require a fee to upload data, but have stricter storage limits.
@@
A final consideration when archiving is the ease of which content can be uploaded by researchers and then downloaded by interested parties. The upload process should be as simple and automated as possible and ideally mirror how researchers already transfer their data between storage options. The Internet Archive, for example, supports uploads of data with simple file management tools like rclone [27], while the Dataverse Project provides ingestion plugins for popular data management tools like iRODS, Globus, and DataLad [48]. In order to support accessible downloads, archives should ideally preserve artifact file structure in URL paths so that deposition files can be cloned via git as discussed in the next section. Stanford’s digital repository exposes raw file structure in this way and the current best public alternative for supporting this feature is the Internet Archive, which unfortunately does not provide immutability or DOI support. Zenodo is overall the best choice for ARTS-compatible deposition at the moment even though it obfuscates individual file paths as a zip archive. A gold standard archive service would also host a container registry in order to centralize and make accessible computational environments used to process deposited data. A full visual comparison of the discussed archive options is provided in Figure 2.
@@ 3.2 Software version control
Git is a “fast, scalable, distributed revision control system” [18] and is the most popular among version control tools which allow researchers to keep a record of experimental code and supporting files [38]. Like the Linux kernel, Git is licensed mostly under the open source GPLv2 license and is available without cost [17]. Git is great for keeping track of code and other lightweight, human readable files, but is generally not advised for use and tracking of large data files.
@@
Git offers the git clone command to copy an entire git tree (files and history) into a new location. Accessing experiment code should be done by running this command and is often done by supplying a link to a git-hosting service such as GitHub. This for-profit service, now owned by Microsoft, does not provide persistent URLs or immutable repositories, but has made a commitment to archiving open source repos [21], has made specific archival efforts [23], and provides integrations for archiving repositories to Zenodo and figshare [22]. That said, repository owners on GitHub may delete the repository or edit its history, breaking published links and calling provenance into question.
Ideally, the code should be directly downloadable from the archive using git clone. This is possible by placing a bare git repo (the .git folder created with git init --bare) at the top-level of a desired repository to share. Most repository web servers will not support the Git HTTP smart service protocol extensions [20], and thus repositories should also be prepared with the git update-server-info command before deposition [19]. This functionality is currently supported by the Internet Archive and SDR.
@@ 3.3 Containers
The backbone of reproducibility in the ARTS open framework is the packaging of the workflow inside a container. Containers offer the promise of executing the same environment on multiple host platforms.
Containers are executed in and managed by container engines. No specific container engine is prescribed, but engines supporting the Open Container Initiative (OCI)-compatible images such as Podman or Docker are recommended. This ensures that as the technology evolves, there will be a way to run the container in the future even if the specific runtime used by the original authors is no longer supported.
Containers are specified by Containerfiles which define any relevant predefined base environment, commands to run and file structures to set up in that environment, and a jumping off point into the environment (entrypoint). Containers can take an environment variables file as an input during runtime—in this case the computational configuration config.env file defined earlier. Files in the local file system can also be made accessible within the container by using volumes—essentially mounting a directory from the local file system within the container.
@@
  • run analysis and generate static research outputs (e.g., figure plots)
  • run interactive development environment (e.g., jupyter notebook server)
  • run interactive command line interface (e.g., ipython kernel)
Ideally, these scripts should be implemented early in the research lifecycle to provide convenience during analysis instead of acting as a barrier to submission. Additional commands provided with the container to perform actions like generating interactive plots and complete manuscript figures, or running data collection pipelines should also be included when possible.

### v2

@@
Sabar Dasgupta Department of Bioengineering Paul Nuyujukian Department of Bioengineering Department of Neurosurgery Department of Electrical Engineering Wu Tsai Neurosciences Institute
@@ 1 Abstract
Digital computational outputs are now ubiquitous in the research workflow and the way in which these data are stored and cataloged is becoming more standardized across fields. However, even with accessible data and code, the barrier to recreating figures and reproducing scientific findings remains high. One element generally missing is the computing environment and associated pipelines in which the data and code are executed to generate figures. The archival, reproducible, and transparent science (ARTS) open framework incorporates containers, version control systems, and persistent archives through which all data, code, and figures related to a research project can be stored together, easily recreated, and serve as an accessible platform for long-term sharing and validation. If the underlying principles behind this framework are broadly adopted, it will improve the reproducibility and transparency of research.
@@ 2 Introduction
Over the past few decades, researchers have recognized the large effort required to replicate experiments and the rate at which these replication studies produce negative results [1]. Dubbed a “replicability crisis” [2], the current paradigm for communicating experimental methods and distributing research outputs impedes future collaboration and the development of public trust in findings. To improve replicability, both the processes for data collection and those for creating reproducible analyses must be improved. Reproducibility is the primary focus of this writing with the goal of minimizing the time, effort, and expertise required to analyze data from a study beyond its published manuscript [3].
@@
By the end of of 2025, research funded by the US federal government must make data accessible at the time of publishing [5, 6]. Each funding agency has their own specific policies for data sharing, but all focus attention on timely release of data in a findable manner [7, 8, 9, 10, 11, 12]. While these policies offer guidance for researchers pursuing collaboration and open science, they do not outline how data, and more importantly, entire workflows should be shared and made accessible. The details around implementing computational best practices are left up to researchers, which is generally outside their scope of expertise and takes the focus away from core scientific work.
There are a number of standards put out by the scientific community to address these concerns ranging from generic (and not necessarily research-specific) to highly focused on one field of research. Data standards such as the FAIR Guiding Principles [13] have gained popularity and are supported by open systems such as DataFed [14]. While FAIR sets a high standard regarding the sharing of data, it alone does not assure reproducibility as it is only starting to be applied to computation [15]. Standards for reproducibility must advise on not only how data is stored, but also how it is manipulated and presented in analysis.
@@
Frameworks such as Spyglass (for neuroscience) or Galaxy (for bioinformatics) recommend field-specific workflows with associated software and are an excellent step towards integrating various software frameworks, but can be fragile as they rely on the permanence and compatibility of external services, software repositories, and APIs [16, 17]. Journals have also recognized the importance software quality and availability best practices [18]. Specifically, both Nature and Science journals now prohibit the use of GitHub for code depositions due to its mutability [19, 20] and Cell strongly recommends archives that support digital longevity [21]. At the time of this writing, even data-focused journals such as Nature Scientific Data provides comprehensive requirements and recommendations for deposition of data and code, but does not prescribe methods for sharing computing environments to promote accessible pipeline reproducibility [22].
@@
It is becoming widely known that consistent computing environments are a crucial piece of scientific reproducibility [23, 24, 25, 26, 27, 28]. Software-as-a-Service (SaaS) tools such as Code Ocean and Brainlife offer reproducibility and analysis pipeline visualization in a standardized environment. [29, 30]. Code Ocean is hosted exclusively via a cloud provider which offers simplicity and reliability at the cost of openness. However, Code Ocean does allow exporting of Containerfiles—the files that define the workflow environment—so that workflow build processes can be run outside of their platform. While Code Ocean is a useful tool for increasing reproducibility for the journal review process [31], it does not define an open standard and is a centralized, commercial solution that does not fully incentivize the sharing of environments throughout the research lifecycle. Brainlife does offer the ability to run jobs on local computing hardware, but is neuroscience-specific. The Whole Tail is yet another promising initiative and stresses the importance of cataloging software environments, but does not appear to have an active community at the time of writing [32]. The calkit Python package also serves as a useful reproducibility tool and acts as a wrapper around some of the technologies discussed later on [33]. With the potential exception of the STARS framework for discrete event simulation healthcare models [34], existing environment-sharing solutions do not yet define an open framework that aims to minimize compliance burden and the time and effort required to reproduce a research workflow many years down the line. Neither highly opinionated field-specific frameworks nor existing generic research workflow aids provide open, end-to-end support at the computational environment level for a wide array of fields.
@@
Heil et al. introduce a medal scale pertaining to computational life science workflows that recommends automating dependency installation and describes a gold standard where analyses are reproducible with a signal command [3]. Unless these dependencies are versioned and running in a consistent environment, there is no guarantee that they will produce deterministic outputs or even run without error in the future. While reproduction effort is a crucial metric, a further key consideration is how far into the future the workflow is able to be run in one step. To truly meet this gold standard the complete software pipelines associated with experiment analyses must be captured as a container image—a binary file that bundles all dependencies, including the operating system, needed to recreate the experiment outputs. A comprehensive source of guidance related to these concepts is available in the Turing Way handbook [35].
@@
Until recently, reproducing entire research workflows was difficult, if not impossible, without locking in to specific computational platforms. Most published depositions do not specify a complete computing environment which leaves the work of setting up a working environment to researchers attempting to reproduce findings. Current approaches also assume that software library repositories such as PyPI for Python or CRAN for R can serve as scientific archives and will continue to exist in their current state decades into the future. Containers, potentially a pillar of scientific computing, offer a solution to environment management that clearly defines entire workflows while being relatively operating system agnostic across local and cloud settings [36]. This writing describes a framework which aims to be a flexible blueprint that is compatible with as many existing research workflows as possible. Specific recommendations are made that encourage reproducibility and define a compatibility layer for those looking to verify results and collaborate. An example deposition is also provided that demonstrates use of the framework in a signal processing context using Python.
3 Results
3.1 Framework Definition
@@
This section defines the Archival, Reproducible, and Transparent Science (ARTS) open framework for accessibly sharing research involving data collection and computational analyses. The goal of this framework is to provide those performing computational analyses that produce research outputs (in this context, digital artifacts) a set of best practices for designing computational research efforts, archiving data and results, sharing said artifacts, and making it simple for others to reproduce and extend results within a consistent analysis environment. The framework can be used collaboratively, incorporating edits and sharing across multiple individuals and groups. Pieces of the ARTS open framework may also be omitted to define best practices for partial workflows—for example, archiving just datasets, archiving code that operates on already-archived datasets, or archiving a standalone computational environment that is used in many experiments.
@@
  • Data and code provenance timestamped (e.g., RFC 3161)
@@
  • Appropriate data and software licenses
The points outlined above and described in detail below should ideally be incorporated early into the research lifecycle where possible. That way, data collection and analysis steps can benefit from automation and archiving the research deposition poses a minimal burden. Further, having a portable deposition that can be uploaded to multiple archives makes research outputs more resilient and having a unified file structure makes outputs more usable.
@@ 3.2 Archives
One goal of archiving experimental results is to make all aspects of analysis available to future experimenters for decades to come: processing data, running models, and generating figures. In order to achieve this, it is vital that the archive service housing the research output be freely accessible as far as possible into the future, or that depositions are able to be archived in a decentralized manner. When choosing an ARTS-compatible archive, there a few key considerations:
Archives must be publicly accessible, meaning that repositories are available for download without paywalls or logins except when necessary to safeguard sensitive information such as Protected Health Information [22]. Accessible archives must also support a variety of open licenses and allow works to be searchable by included metadata like keywords.
@@
Archives must allow for persistent storage of artifacts (e.g., files and directories). Once results are published, they should remain immutable (i.e., cannot be quietly changed) and allowances for future updates or edits must result in an immutable history that includes the original publication. If history is not immutable, trusted timestamping information must be provided (an example is described below). Persistent archives must also support persistent identifiers of some kind. Ideally this should be a digital object identifier (DOI), but a persistent URL (PURL) that is not a DOI may suffice. If the archive does not support DOIs, published manuscripts describing the data should be minted with a DOI and include relevant PURLs. Centralized persistent archive services should also store a copy of data across multiple geographic regions so that they are resilient to external events that might result in data loss such as natural disasters or political events.
@@
Archives must also be trusted. Requirements for trust cannot be singularly defined in a broad sense, but at a minimum, an archive service should be established with long-term financial support and a dedicated community. Archives backed by noncommercial institutions are also preferable as they are less likely to change their policies due to economic pressures. Providing for an optional confidential review process of artifacts also promotes trust in the archive and the data it houses [22].
@@
Figure 2: No popular research archive currently supports all aspects of the ARTS open framework, but Zenodo supports all aspects outside of hosting a container registry and allowing direct git clones of analysis code. Each column (besides storage limits) defines a question with checks and X’s indicating yes and no, respectively. Columns definitions are as follows: 3rd party upload: Does the archive allow anyone to upload data? Immutable: Once published, is deposition content locked from edits without provenance information? DOI: Does the archive supply a Digital Object Identifier for uploads? PURL: Does the archive supply a persistent URL for uploads (this may be a DOI)? Geographic redundancy: Is data uploaded to the archive backed up in separate geographic locations? Reviews: Does the archive provide a process for privately and confidentially sharing uploads to reviewers before publishing? Nonprofit: Does the maintainer of the archive operate under nonprofit status? Storage limits: Maximum file and deposition size as well as any other deposition limits. Bare repo: Does the archive allow depositions to be directly git cloned? Container pull: Does the archive implement a container registry that allows direct image pulls from container engines?
@@
Some academic institutions offer in-house archive solutions: the Dataverse Project originally developed at Harvard [37, 38], DSpace developed between MIT and HP [39], and the Stanford Digital Repository (SDR) [40] are some examples of archives that promote open access and provide persistent identifiers. DSpace and Dataverse are open-source tools with instances hosted by many partner organizations while the SDR is only available to Stanford affiliates. There are also popular and journal-recommended third-party solutions such as Zenodo [41], Dryad [42], or figshare [43] if a field-specific data repository cannot be found. Dryad allows up to 300GB of storage per submission with a publication fee, while Zenodo and figshare do not require a fee to upload data, but have stricter storage limits.
@@
A final consideration when archiving is the ease by which content can be uploaded by researchers and then downloaded by interested parties. The upload process should be as simple and automated as possible and ideally mirror how researchers already transfer their data between storage options. The Internet Archive, for example, supports uploads of data with simple file management tools like rclone [44], while the Dataverse Project provides ingestion plugins for popular data management tools like iRODS, Globus, and DataLad [45]. Allowing researchers to mint DOIs before upload so they may include them in the first version of their deposition is also helpful. In order to support accessible downloads, archives should ideally preserve artifact file structure via a static URL path and common URL prefix for the entire deposition so that files can be cloned via git as discussed in the next section. Stanford’s Digital Repository exposes raw file structure in this way and the current best public alternative for supporting this feature is the Internet Archive, which unfortunately does not provide immutability or DOI support. Zenodo is overall the best choice for ARTS-compatible deposition at the moment even though it obfuscates individual file paths as a zip archive. A gold standard archive service would also host a container registry in order to centralize and make accessible computational environments used to process deposited data. A full visual comparison of the discussed archive options is provided in Figure 2.
@@ 3.3 Software version control
Git is a “fast, scalable, distributed revision control system” [46] and is the most popular among version control tools which allow researchers to keep a record of experimental code and supporting files [36]. Like the Linux kernel, Git is licensed mostly under the open source GPLv2 license and is available without cost [47]. Git excels at keeping track of code and other lightweight, human readable files, but is generally not advised for use and tracking of large data files.
@@
Git offers the git clone command to copy an entire git tree (files and history) into a new location. Accessing experiment code should be done by running this command and is often done by supplying a link to a git-hosting service such as GitHub. This for-profit service, now owned by Microsoft, does not provide persistent URLs or immutable repositories, but has made a commitment to archiving open source repositories [48], has taken specific actions to this effect [49], and provides integrations for archiving repositories to Zenodo and figshare [50]. That said, repository owners on GitHub may delete the repository or edit its history, breaking published links and calling provenance into question.
Ideally, code should be directly downloadable from the archive using git clone. This is possible by placing a bare git repository (the .git folder created with git init --bare) at the top-level of a desired repository to share. Most repository web servers will not support the Git HTTP smart service protocol extensions [51], and thus repositories should also be prepared with the git update-server-info command before deposition [52]. This functionality is currently supported by the Internet Archive and the Stanford Digital Repository.
@@ 3.4 Containers
The backbone of reproducibility in the ARTS open framework is the packaging of the workflow inside a container. Containers offer the promise of executing the same environment on multiple host platforms and are used by services such as Binder to provide a shared computing environment [53]. Their use is growing across scientific fields especially to drive collaborative and reproducible efforts [54].
Containers are executed in and managed by container engines. No specific container engine is prescribed by ARTS, but engines supporting Open Container Initiative (OCI)-compatible images such as Podman or Docker are recommended [55]. This ensures that as the technology evolves, there will be a way to run the container in the future even if the specific runtime used by the original authors is no longer supported.
Containers are specified by Containerfiles which define any relevant predefined base environment, commands to run and file structures to set up in that environment, and a jumping off point (i.e., default execution script) into the environment (entrypoint). Containers can take environment variables as inputs during runtime—in this case the computational configuration config.env file defined below. Files in the local file system can also be made accessible within the container by using volumes to mount a directory from the local file system within the container.
@@
  • run analysis and generate static research outputs (e.g., figures or numerical results)
  • run interactive development environment (e.g., jupyter server)
  • run interactive command line interface (e.g., IPython kernel or IRkernel)
Ideally, these scripts should be implemented early in the research lifecycle to provide convenience during the exploratory and iterative phases of research instead of acting as a barrier to submission. Additional commands provided with the container to perform actions like generating interactive plots and complete manuscript figures, or running data collection pipelines should also be included when possible.